Find the product. 
(a^2+ b) (a + b^2 )
Find the product. 
(a^2+ b) (a + b^2 )

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(a^2+ b) (a + b^2)
= a^2(a + b^2) b(a + b^2)
= a^3+ a^2b^2 + ab + b^3
= a^3+ b^3+ a^2b^2 + ab

Thanks